HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

DIFFERENT NON-COVALENT INTERACTION TYPES AND GEOMETRIES FOR SIMILAR COMPLEXES OF

PHENYLACETALDEHYDE–H_{2}O AND PHENYLACETALDEHYDE–H_{2}S

\underline{M. LI,} W. LI, J.-U. GRABOW, A. LESARRI, Institut für Physikalische Chemie und Elektrochemie, Leibniz Universität Hannover, Hannover 30167, (Germany); M.

Li: meng.li@pci.uni-hannover.de

Molecules with aromatic rings are extremely versatile in exercising non-covalent interactions (NCIs), making them prototype systems to study π–π stacking,^{1} lone pair...π–hole interactions,^{2} and H...π interactions,^{3} etc. Phenylacetaldehyde (PAA) is the derivative of the simplest aromatic aldehyde (benzaldehyde) towards a longer carbon chain. Thus, there are more probabilities for its bonding with other molecules by different NCI sites due to its increased flexibility.

The chalcogens O and S, being homologues next to each other in their main group of the periodic system, show some similar but indeed varying physical and chemical properties. Therefore, H_{2}O and H_{2}S can serve as simple model molecules to study how such property changes effect NCI types and, consequently, the geometries of their respective complexes formed between them and identical partner molecules.

E.g., it had been reported that the geometry is similar in the case of benzene– H_{2}O^{4} and benzene–H_{2}S complexes,^{5} but very different in phenylacetylene–H_{2}O^{6} and phenylacetylene–H_{2}S complexes.^{7} This raises the question if trends for different functionalizatin can be established: E.g., how will the aldehyde complexes of PAA– H_{2}O and PAA–H_{2}S behave? At this purpose, the two title complexes have been investigated using broadband rotational spectroscopy in frequency range of 2–8 GHz.

1doi:acs.jpclett.2c02807, W. Du, Y. Zheng, X. Wang, J. Lei, H. Wang, X. Tian, S. Zou, J. Bloino, Q.

Gou, W. Caminati and J.–U. Grabow, J. Phys. Chem. Lett., 13 (42), 9934–9940 (2022).

2doi:acs.jpclett.6b00473, C. Calabrese, Q. Gou, A. Maris, W. Caminati and S. Melandri, J. Phys.

Chem. Lett., 7 (8), 1513–1517 (2016).

3doi:jacs.2c11732, W. Li, C. Pérez, A. L. Steber, M. Schnell, D. Lv, G. Wang, X. Zeng, and M.

Zhou, J. Am. Chem. Soc., 145 (7), 4119–4128 (2023).

4doi:acs.jpca.9b01497, J. Andersen, R. W. Larsen, J. Ceponkus, P. Uvdal and B. Nelander, J. Phys.

Chem. A., 124 (3), 513–519 (2019).

5doi:acs.jpca.9b02199, D. Wang, P. Chopra, S. Wategaonkar and A. Fujii, J. Am. Chem. Soc., 145 (7), 4119–4128 (2023).

6doi:C1CP20690G, M. Goswami and E. Arunan, Phys. Chem. Chem. Phys., 13 (31), 14153–14162 (2011).

7doi:j.jms.2011.04.011, 8. M. Goswami and E. Arunan, J. Mol. Spectrosc., 268 (1–2), 147–156 (2011).

p-number: p040  Submitted on Tue, 06 Jun 2023 14:09:32 +0200